Chapter 15, Problem 10 l Your answer is partially correct. Try again Two tiny conducting spheres are identical and carry charges of -19.7 C and +68 4HC They are separated by a distance of 241 cm. (a) what is the magnitude of t each sphere experiences? (b) The spheres are brought into contact and then separated to a distance of 2.41 cm, Determine the magnitude of the force that each sphere now experiences (a) N (b) Number
Answer #1

Case 1

1F 19.TuC=-1.97 * 10-5C

68.4uC 6.84 10

d 2.41cm 0.0241771

kq142

F 9 11.97 106.84 10 /(0.0241m)-

F 9* 101.97 106.84100.0241

F=12.127/0.0241^{2}

ANSWER: {color{Red} F=20880N} (attractive)

=============

When the spheres are brought in contact ,the charges equalizes

68.4-19.7- 48.7

this charge will be equally distributed

gi = 92 = 24.35μ C-2.435 * 10-5C

d 2.41cm 0.0241771

kq142

F=9*10^{9}*2.435*10^{-5}C*2.435*10^{-5}C/(0.0241m)^{2}

F=9*10^{9}*2.435*10^{-5}*2.435*10^{-5}/0.0241^{2}

F 5.336/0.0241 ,330

ANSWER: {color{Red} F=9187.69N} (repulsive)

================================